OpenStudy (anonymous):

Use a transformation to solve the equation. d – (–17) = –14 A. d = –31 B. d = –3 C. d = 3 D. d = 31

OpenStudy (alexandervonhumboldt2):

add to both sides -17

OpenStudy (alexandervonhumboldt2):

you will get d-(-17)+(-17)=-14+(-17 this is equal to d=-14-17=?

OpenStudy (anonymous):

so the answer is -31?

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (whpalmer4):

Put -31 in the original equation and check it to see! \[-31 - (-17) = -14\] \[-31 + 17 = -14\] \[-31 +17 - 17 = -14-17\] \[-31 = -31\checkmark\] You have a true statement when you put the proposed answer in the original equation, therefore it is correct.

OpenStudy (anonymous):

oh, ok. thank you @whpalmer4 @AlexandervonHumboldt2